OpenStudy (anonymous):

find the general solution by separating the variables: x(1+y^2)dx +y(1+x^2)dy=0

OpenStudy (anonymous):

x(1 + y^2) = -y(1 + x^2)dy

OpenStudy (anonymous):

xdx / (1 + x^2) = -ydy/(1 + y^2)

OpenStudy (anonymous):

For the left hand side to take the integral let u = 1 + x^2 then du = 2xdx so du/2 = xdx. Also for the right hand side let u = 1+ y^2 then du = 2ydy which also means that -du/2 = -ydy

OpenStudy (anonymous):

So know you have

OpenStudy (anonymous):

\[1/2\int\limits_{}^{} du/u = -1/2\int\limits_{}^{} du/u\]

OpenStudy (anonymous):

1/2ln|u| = -1/2lln|u| ln|1+ x^2| = -ln|1 + y^2] + C

OpenStudy (anonymous):

Then you take e to the power of both sides: getting rid of the natural logs: 1+ x^2 = e^c(1 + y^2)...keep in mind that e^c is another constant lets say K.

OpenStudy (anonymous):

1/k(1 + x^2) = 1 + y^2 y^2 = 1 - 1/k(1 + x^2) y = sqrt(1 - 1/k(1 +x^2))

OpenStudy (anonymous):

It might also be clearer and more understandable if for the integration you set different variables for your denominators instead of using u for both the left and right hand side.

OpenStudy (anonymous):

do you follow?

OpenStudy (anonymous):

yea, the problem I had was that I messed up after the separation and made a large mess out of the integration. Everything after the integration is the easier part for me.

OpenStudy (anonymous):

Or I should say finding the proper integration method was my problem.

OpenStudy (anonymous):

okay yeah with separation of variables dividing and setting up both sides to integrate is key. It's good that you understand though but I assure you that as you do more practice with separation of variables you will master various integration techniques as well.

OpenStudy (anonymous):

after I write this one down i'll be trying out a few more and hopefully that is the case since practice should make perfect :)

OpenStudy (anonymous):

exactly! that's how I learnt it! but if you have any questions don't hesitate to post them.

OpenStudy (anonymous):

okay, thanks alot for helping me out

OpenStudy (anonymous):

no problem